Cardinal Industries had the following operating results for 2018: Sales $34,722; Cost of goods sold $24,408; Depreciation expense $6,037; Interest expense $2,730; Dividends paid $2,032. At the beginning of the year, net fixed assets were $19,980 current assets were $7,082, and current liabilities were $4,016. At the end of the year, net fixed assets were $24,538, current assets were $8,708, and current liabilities were $4,709. The tax rate for 2018 was 21 percent. a. What is net income for 2018? (Do not round intermediate calculations.) b. What is the operating cash flow for 2018? (Do not round intermediate calculations.) c. What is the cash flow from assets for 2018? (Do not round intermediate calculations. A negative answer should be indicated by a minus sign.) d- If no new debt was issued during the year, what is the cash flow to creditors? (Do not 1. round intermediate calculations.) d- If no new debt was issued during the year, what is the cash flow to stockholders? (Do 2. not round intermediate calculations. A negative answer should be indicated by a minus sign.) a. Net income b. Operating cash flow c. Cash flow from assets d-1. Cash flow to creditors d-2. Cash flow to stockholders 1,222
Answer #1

FY2018 FY2017 FY2018   

Sales 34722 net fixed assets 19980 24538

less cost of good sold -24408 current assets 7082 8708

depriciation - 6037 current liabilities 4016 4709

intrest -2730 working capital 3066 3999

pbt 1547 working capital = current assets - current liabilities

tax 0.21*1547 =309..40

Q a) net income 1547 - 309.40= 1237.60

dividends 2032

Q B) operating cash flow = net income +depriciation+increase in working capital

operating cash flow = 1237..60+ 6037( 3999- 3066)

= 6341

Q C) Cash flow from assets = capex = --(increased in net fixed assets+ depriciation)

cash flow from assets = --10595

Q D) Cash flow to share holders = operating cf -capex + net debt issued

cash flow to shareholders = 6341 + (-10595)

= 4254